# Neural Topic Modeling with Large Language Models in the Loop

### v2

@@ Appendix A Prompts
Figure A1: Prompt and output of topic suggestion with CoT. We take the product of token probabilities of topic label (e.g., words in red color) as the label token probability. We take the proportion of intruders (e.g., words in green color) as the word intrusion confidence. $N=2$ and $M=10$ in this example.
@@ Appendix B Algorithm
Input: Train documents; An LLM; Pre-trained word embeddings; Hyper-parameters $T^{\text{refine}}$, $\gamma$; Training iteration $I$; Number of topics $K$.
@@ Appendix C Related work: LLM Uncertainty Estimation
Uncertainty estimation for LLMs (Geng et al., 2024) is emerging with the rapid usage of LLMs and their risk of hallucination (Ji et al., 2023). Sequence probability (Ren et al., 2022) leverages token probabilities to measure answer confidence. Verbalized confidence (Tian et al., 2023; Xiong et al., 2023) utilizes the LLM’s own capability to evaluate its answer uncertainty. Consistency-based (Lin et al., 2023; Manakul et al., 2023) approaches sample multiple outputs from the LLM and measure answer consistency as uncertainty. Entropy-based (Kuhn et al., 2023; Hou et al., 2023) approaches estimate the output space from multiple LLM outputs and compute entropy as uncertainty for the answer. Hybrid frameworks (Chen and Mueller, 2023; Gao et al., 2024) combine different approaches for a comprehensive estimation. Internal states (Chen et al., 2024) are another useful source for LLM uncertainty quantification. Unlike those works that estimate uncertainty for LLMs in general natural language generation tasks, ours focuses on task-specific uncertainty of the LLM in suggesting topic words.
@@ D.1 Details of Dataset
We conduct experiments on 20News [^3] https://huggingface.co/datasets/SetFit/20_newsgroups, R8444https://huggingface.co/datasets/yangwang825/reuters-21578, DBpedia [^5] https://huggingface.co/datasets/fancyzhx/dbpedia_14, and AGNews [^6] https://huggingface.co/datasets/fancyzhx/ag_news. For DBpedia and AGNews, we randomly sample a subset of 20,000 documents. We retain the original text documents for models that accept text as input, and preprocess the documents into Bag-of-Words (BOW) format for models that are trained on BOWs. We convert the documents into BOW vectors through the following steps: First, we clean the documents by removing special characters and stop words, followed by tokenization. Next, we build the vocabulary by including words with a document frequency greater than five and less than 80% of the total documents. Since we use the pre-trained word embeddings of GloVe (Pennington et al., 2014), we further filter the vocabulary by retaining only the words that are in the GloVe vocabulary. Finally, we transform the documents into BOWs based on the filtered vocabulary set. The statistics of the preprocessed datasets are summarized in Table D1.
@@ D.2 Details of Baselines
We run the following topic models in our experiments, including Latent Dirichlet Allocation (LDA) (Blei et al., 2003), the most popular probabilistic topic model that generates documents by mixtures of topics; Neural Variational Document Model (NVDM) (Miao et al., 2017), a pioneering NTM based on the VAE framework; LDA with Products of Experts (PLDA) (Srivastava and Sutton, 2017), an NTM that uses a product of experts instead of the mixture model in LDA; Neural Topic Model with Covariates, Supervision, and Sparsity (SCHOLAR) (Card et al., 2017), an NTM that leverages extra information from metadata; Embedded Topic Model (ETM) (Dieng et al., 2020), which involves word and topic embeddings in the generative process of documents; Neural Sinkhorn Topic Model (NSTM) (Zhao et al., 2020), an NTM based on an optimal transport framework; Contrastive Learning Neural Topic Model (CLNTM) (Nguyen and Luu, 2021), an NTM that is based on the contrastive learning framework; BERTopic (Grootendorst, 2022), a recent clustering-based topic model that utilizes embeddings from pre-trained language models; WeTe (Wang et al., 2022), which represents a mixture of word embeddings as a mixture of topic embeddings; Embedding Clustering Regularization Topic Model (ECRTM) (Wu et al., 2023), an NTM with embedding clustering regularization; TopicGPT (Pham et al., 2023), a latest LLM-based topic model that leverages an LLM for topic generation and assignment.
@@ D.3 Details of Evaluation Metrics
For topic evaluation, we apply the commonly-used topic coherence metric, $\bm{C_{V}}$ (Röder et al., 2015), which evaluates topic coherence based on the co-occurrence of the topic’s top words in a reference corpus. Following standard evaluation protocol, we use Wikipedia as the reference corpus and consider the top 10 words of each topic, with implementation done using the Palmetto package [^9] https://github.com/dice-group/Palmetto (Röder et al., 2015). We report the average $C_{V}$ score of all learned topics as the overall topic coherence performance. For documents’ topical representation (i.e., topic proportion) evaluation, a common practice is to compare the document clusters formed by topic proportions with those formed by the documents’ true labels, known as topic alignment. Following previous works (Chuang et al., 2013; Pham et al., 2023), we assign each test document to a cluster based on the top-weighted topic of its topical representation, and compute Purity and Normalized Mutual Information (NMI) based on the documents’ cluster assignments and their true labels. As Purity and NMI are often reported together and within the same range, we report the average score of both metrics as PN. For all evaluations, we use the model state at the end of the training iteration to compute all evaluation metrics.
@@ Appendix F Study of Confidence Alternatives
Here, we compare our proposed topic labeling confidence with other LLM confidence alternatives within the LLM-ITL framework. We limit our focus to single-sample approaches, where a single round of LLM inference is performed for a given topic. This ensures they can be efficiently applied within the training of LLM-ITL. We consider the following confidence alternatives to our topic labeling confidence in this study: (1) No Conf., where no LLM confidence estimation is included, and $\text{Conf}(\bm{w}_{k}^{l})=1$ in Eq. 10 during the training. (2) Verbalized Confidence (Xiong et al., 2023), which directly asks the LLM for its confidence in solving a problem. The prompt we used for eliciting verbalized confidence is shown in Figure A2. (3) Self-Reflective Confidence (Chen and Mueller, 2023), which prompts the LLM to evaluate its own answer in a two-stage manner. The topic label is obtained in the first stage, and the LLM evaluates this answer in a follow-up question (Figure A3). (4) p(True) (Kadavath et al., 2022), which is similar to self-reflective confidence, but asks a true/false question instead. It takes the token probability of the response as the confidence measure (Figure A3). (5) SeqLike. (Ren et al., 2022), which computes the length-normalized sequence likelihood of the output from the LLM.
@@ Appendix H Study of Prompts
Here, we study the effectiveness of different prompts for topic suggestion. We obtain variants of topic suggestion prompts in Figure A1 through modifying the technique used in PromptBreeder (Fernando et al., 2023). To be specific, we first create a set of 100 “mutation-prompts” (e.g., “Make a variant of the prompt") and 100 “thinking-styles” (e.g., “Let’s think step by step"). We generate a set of 50 task prompts by concatenating a randomly drawn “mutation-prompt” and a randomly drawn “thinking-style” to the original prompt, and provide that to the Claude 3.5101010https://www.anthropic.com/news/claude-3-family to produce a continuation, resulting in a different task prompt. Secondly, we randomly select 50 topics over 4 experiment datasets. We run those topics through 50 generated task prompts and filter out the generated prompts that cannot give JSON format in the selected topics or generate above 300 tokens. We are left with 14 topics. We then leverage Claude 3.5 to judge the quality of generated topics and refined topic words. We rank 14 methods by overall topics and refined topic words to get 5 variants of prompts. In addition to the prompt variants generated by those steps, we also investigate the topic refinement prompt used in Chang et al. (2024) (see Figure 2 of their paper). All the prompt variants for topic refinement in this study are illustrated in Table H2.

### v3

@@ 5.2 Results
Ablation Study of OT
$\bm{C_{V}}$ PN
KL 0.480 ± 0.007 0.402 ± 0.012
JSD 0.479 ± 0.003 0.403 ± 0.011
HD 0.480 ± 0.011 0.403 ± 0.012
TVD 0.480 ± 0.008 0.405 ± 0.011
OT 0.578 ± 0.008 0.398 ± 0.010
Table 3: Ablation study of OT. The best performance is highlighted in boldface for both metrics.
Here, we conducted an ablation study on our OT-based topic refinement approach. In this study, we apply different metrics to measure the differences between the topic word distributions generated by the NTM and those refined by the LLM. Specifically, we explored alternative metrics to optimal transport (OT) within the LLM-ITL framework, including Kullback–Leibler (KL) divergence, Jensen–Shannon divergence (JSD), Hellinger distance (HD), and Total Variation distance (TVD).
As shown in Table 3, which presents results from experiments on 20News with $K=50$ using ETM as the base model, our OT-based approach demonstrates a significant advantage in improving topic coherence compared to alternative distribution measurement methods.
@@ 7 Limitation
Acknowledgment
We thank the anonymous (meta) reviewers in ARR for their valuable feedback, which has significantly strengthened this work.
@@ Appendix B Algorithm
Input: Train documents; An LLM; pre-training word embeddings; Hyper-parameters $T^{\text{refine}}$, $\gamma$; Training iteration $I$; Number of topics $K$.
@@ Appendix C Related work: LLM Uncertainty Estimation
Uncertainty estimation for LLMs (Geng et al., 2024) is emerging with the rapid usage of LLMs and their risk of hallucination (Ji et al., 2023). Sequence probability (Ren et al., 2023) leverages token probabilities to measure answer confidence. Verbalized confidence (Tian et al., 2023; Xiong et al., 2024) utilizes the LLM’s own capability to evaluate its answer uncertainty. Consistency-based (Lin et al., 2024; Manakul et al., 2023) approaches sample multiple outputs from the LLM and measure answer consistency as uncertainty. Entropy-based (Kuhn et al., 2023; Hou et al., 2024) approaches estimate the output space from multiple LLM outputs and compute entropy as uncertainty for the answer. Hybrid frameworks (Chen and Mueller, 2024; Gao et al., 2024b) combine different approaches for a comprehensive estimation. Internal states (Chen et al., 2024) are another useful source for LLM uncertainty quantification. Unlike those works that estimate uncertainty for LLMs in general natural language generation tasks, ours focuses on task-specific uncertainty of the LLM in suggesting topic words.
@@ D.1 Details of Dataset
We conduct experiments on 20News [^3] https://huggingface.co/datasets/SetFit/20_newsgroups, R8444https://huggingface.co/datasets/yangwang825/reuters-21578, DBpedia [^5] https://huggingface.co/datasets/fancyzhx/dbpedia_14, and AGNews [^6] https://huggingface.co/datasets/fancyzhx/ag_news. For DBpedia and AGNews, we randomly sample a subset of 20,000 documents. We retain the original text documents for models that accept text as input, and preprocess the documents into Bag-of-Words (BoW) format for models that are trained on BoWs. We convert the documents into BoW vectors through the following steps: First, we clean the documents by removing special characters and stop words, followed by tokenization. Next, we build the vocabulary by including words with a document frequency greater than five and less than 80% of the total documents. Since we use the pre-training word embeddings of GloVe (Pennington et al., 2014), we further filter the vocabulary by retaining only the words that are in the GloVe vocabulary. Finally, we transform the documents into BoWs based on the filtered vocabulary set. The statistics of the preprocessed datasets are summarized in Table D1.
@@ D.2 Details of Baselines
We run the following topic models in our experiments, including Latent Dirichlet Allocation (LDA) (Blei et al., 2003), the most popular probabilistic topic model that generates documents by mixtures of topics; Neural Variational Document Model (NVDM) (Miao et al., 2017), a pioneering NTM based on the VAE framework; LDA with Products of Experts (PLDA) (Srivastava and Sutton, 2017), an NTM that uses a product of experts instead of the mixture model in LDA; Neural Topic Model with Covariates, Supervision, and Sparsity (SCHOLAR) (Card et al., 2018), an NTM that leverages extra information from metadata; Embedded Topic Model (ETM) (Dieng et al., 2020), which involves word and topic embeddings in the generative process of documents; Neural Sinkhorn Topic Model (NSTM) (Zhao et al., 2021b), an NTM based on an optimal transport framework; Contrastive Learning Neural Topic Model (CLNTM) (Nguyen and Tuan, 2021), an NTM that is based on the contrastive learning framework; BERTopic (Grootendorst, 2022), a recent clustering-based topic model that utilizes embeddings from pre-trained language models; WeTe (dongsheng wang et al., 2022), which represents a mixture of word embeddings as a mixture of topic embeddings; Embedding Clustering Regularization Topic Model (ECRTM) (Wu et al., 2023), an NTM with embedding clustering regularization; TopicGPT (Pham et al., 2024), a latest LLM-based topic model that leverages an LLM for topic generation and assignment.
@@ D.3 Details of Evaluation Metrics
For topic evaluation, we apply the commonly-used topic coherence metric, $\bm{C_{V}}$ (Röder et al., 2015), which evaluates topic coherence based on the co-occurrence of the topic’s top words in a reference corpus. Following standard evaluation protocol, we use Wikipedia as the reference corpus and consider the top 10 words of each topic, with implementation done using the Palmetto package [^9] https://github.com/dice-group/Palmetto (Röder et al., 2015). We report the average $C_{V}$ score of all learned topics as the overall topic coherence performance. For documents’ topical representation (i.e., topic proportion) evaluation, a common practice is to compare the document clusters formed by topic proportions with those formed by the documents’ true labels, known as topic alignment. Following previous works (Chuang et al., 2013; Pham et al., 2024), we assign each test document to a cluster based on the top-weighted topic of its topical representation, and compute Purity and Normalized Mutual Information (NMI) based on the documents’ cluster assignments and their true labels. As Purity and NMI are often reported together and within the same range, we report the average score of both metrics as PN. For all evaluations, we use the model state at the end of the training iteration to compute all evaluation metrics.
@@ Appendix F Study of Confidence Alternatives
Here, we compare our proposed topic labeling confidence with other LLM confidence alternatives within the LLM-ITL framework. We limit our focus to single-sample approaches, where a single round of LLM inference is performed for a given topic. This ensures they can be efficiently applied within the training of LLM-ITL. We consider the following confidence alternatives to our topic labeling confidence in this study: (1) No Conf., where no LLM confidence estimation is included, and $\text{Conf}(\bm{w}_{k}^{l})=1$ in Eq. 10 during the training. (2) Verbalized Confidence (Xiong et al., 2024), which directly asks the LLM for its confidence in solving a problem. The prompt we used for eliciting verbalized confidence is shown in Figure A1. (3) Self-Reflective Confidence (Chen and Mueller, 2024), which prompts the LLM to evaluate its own answer in a two-stage manner. The topic label is obtained in the first stage, and the LLM evaluates this answer in a follow-up question (Figure A2). (4) p(True) (Kadavath et al., 2022), which is similar to self-reflective confidence, but asks a true/false question instead. It takes the token probability of the response as the confidence measure (Figure A2). (5) SeqLike. (Ren et al., 2023), which computes the length-normalized sequence likelihood of the output from the LLM.
@@ Appendix H Study of Prompts
Here, we study the effectiveness of different prompts for topic suggestion. We obtain variants of topic suggestion prompts in Figure 2 through modifying the technique used in PromptBreeder (Fernando et al., 2024). To be specific, we first create a set of 100 “mutation-prompts” (e.g., “Make a variant of the prompt") and 100 “thinking-styles” (e.g., “Let’s think step by step"). We generate a set of 50 task prompts by concatenating a randomly drawn “mutation-prompt” and a randomly drawn “thinking-style” to the original prompt, and provide that to the Claude 3.5101010https://www.anthropic.com/news/claude-3-family to produce a continuation, resulting in a different task prompt. Secondly, we randomly select 50 topics over 4 experiment datasets. We run those topics through 50 generated task prompts and filter out the generated prompts that cannot give JSON format in the selected topics or generate above 300 tokens. We are left with 14 topics. We then leverage Claude 3.5 to judge the quality of generated topics and refined topic words. We rank 14 methods by overall topics and refined topic words to get 5 variants of prompts. In addition to the prompt variants generated by those steps, we also investigate the topic refinement prompt used in Chang et al. (2024) (see Figure 2 of their paper). All the prompt variants for topic refinement in this study are illustrated in Table H2.
@@ Results
Appendix I Topic Quality Based on Distributed Word Representations
Dataset Method W2V-Cosine ()($\downarrow$)( ↓ ) W2V-L1 ()($\downarrow$)( ↓ ) W2V-L2 ()($\downarrow$)( ↓ )
20News ETM 0.218 ± 0.002 2.313 ± 0.017 11.459 ± 0.180
+ LLM-ITL 0.177 ± 0.005 2.128 ± 0.040 9.650 ± 0.335
$\uparrow$ 18.8% $\uparrow$ 8.0% $\uparrow$ 15.8%
R8 ETM 0.237 ± 0.004 2.543 ± 0.023 13.673 ± 0.211
+ LLM-ITL 0.180 ± 0.006 2.245 ± 0.042 10.855 ± 0.345
$\uparrow$ 24.1% $\uparrow$ 11.7% $\uparrow$ 20.6%
DBpedia ETM 0.213 ± 0.003 2.448 ± 0.020 12.709 ± 0.237
+ LLM-ITL 0.163 ± 0.010 2.119 ± 0.090 9.633 ± 0.696
$\uparrow$ 23.5% $\uparrow$ 13.4% $\uparrow$ 24.2%
AGNews ETM 0.203 ± 0.005 2.419 ± 0.023 12.382 ± 0.234
+ LLM-ITL 0.164 ± 0.004 2.175 ± 0.035 10.031 ± 0.241
$\uparrow$ 19.2% $\uparrow$ 10.1% $\uparrow$ 19.0%
Table I1: Topic quality in terms of Word2vec metrics.
In this section, we conduct a performance comparison of topic quality using metrics based on distributed word representations (Nikolenko, 2016), which have been shown to align better with human judgment than standard topic coherence metrics.
Specifically, we use the topic quality metrics based on word vectors as implemented by Nikolenko (2016), applying cosine distance (W2V-Cosine), L1 distance (W2V-L1), and L2 distance (W2V-L2), respectively, as the distance functions in the calculation. As shown in Table I1, LLM-ITL consistently improves semantic coherence of topics, as measured by the Word2Vec-based evaluation metrics.